Question 2A:

Quantity TC TVC TFC ATC AVC MC
0 30
1 40
2 50
3 30
4 30
5 40
6 260
7 300

The respective cost values when the firm's output is 7, are:

360; 300; 30; 57.1; 42.9; 100.
330; 300; 30; 33; 30; 70.
330; 300; 30; 47.1; 42.9; 70.

260; 300; 30; 43.3; 42.9; 0.
Answer #1

Ans: 330; 300; 30; 47.1; 42.9; 70.

Explanation:

The firms TFC is 30. Because when output is 0, TC = 30. We know TC = TFC + TVC. So, when TC = 30 and TVC = 0 at output level of 0. So, TFC = 30

Quantity TC TVC TFC ATC = TC/Q AVC = TVC/Q MC = ∆TC / ∆Q
0 30 0 30 - 0 -
1 70 40 30 70.0 40.0 40
2 100 70 30 50.0 35.0 30
3 120 90 30 40.0 30.0 20
4 150 120 30 37.5 30.0 30
5 200 170 30 40.0 34.0 50
6 260 230 30 43.3 38.3 60
7 330 300 30 47.1 42.9 70